Host A wants to transfer 5000 bytes of data to Host B. On the network path there are two routers. First router R1, has MTU(Maximum Transformation Unit) of 1900 bytes. Second router R2, has MTU of 900 bytes.

  1. How many packets will Router 1 will deliver to Router 2 ?
  2. Draw packet formats with More Flag and Offset values and data sizes for each datagram for R1 .
  3. How many packets will Router 2 will deliver to Host B ?
  4. Draw packet formats with More Flag and Offset values and data sizes for each datagram for R2 sending to Host B.
Answer #1

Answer

1. In Case of 1st router it has MTU of 1900B and data is 5000B>1900B so fragmentation is needed.
So 20B of header info will be added to each packed being sent.
So in one packed we can send 1900 - 20 = 1880B of data,
So we will need 5000/1880 = 2 full packets of 1880B data and one packet having data of 1240B
So Router 1 will send 3 packets.

  • 2 packet of 1880B of data and 20B header
  • 1 packet of 1240B data and 20B header.

2. Packet 1: More_Flag = true, offset_value = 0, data_size = 1880B
   Packet 2: More_Flag = true, offset_value = 1881, data_size = 1880B
   Packet 3: More_Flag = false, offset_value = 3761, data_size = 1240B

3. So as router B has MTU of 900B so it can send 900-20B = 880B of data, so fragmentation is needed for the packets    receiving from router A.
So it will need

  • 1880/880 = 3 packets(2 full 1 partial ) for the first packet of router 1
  • 1880/880 = 3 packets(2 full 1 partial ) for the second packet of router 1
  • 1240/880 = 2 packets(1 full 1 partial ) for the third packet of router 1

So total of 8 packets

4. Following would be packets of router 2.

  1. Packet 1: more_flag = true, offset_value = 0, data_size = 880
  2. Packet 2: more_flag = true, offset_value = 881, data_size = 880
  3. Packet 3: more_flag = true, offset_value = 1761, data_size = 120
  4. Packet 4: more_flag = true, offset_value = 1881, data_size = 880
  5. Packet 5: more_flag = true, offset_value = 2761, data_size = 880
  6. Packet 6: more_flag = true, offset_value = 3641, data_size = 120
  7. Packet 7: more_flag = true, offset_value = 3761, data_size = 880
  8. Packet 8: more_flag = false, offset_value = 4641, data_size = 360